OpenStudy (anonymous):

use the geometric mean to find the 7th term in a geometric if the 6th term is 6 and the 8th is 216

OpenStudy (anonymous):

do you know what "geometric mean" means?

OpenStudy (anonymous):

nope im takin a timed quiz i didnt know how to solve it

OpenStudy (anonymous):

the geometric mean of \(a\) and \(b\) is \(\sqrt{a\times b}\) so your job is to compute \[\sqrt{6\times 216}\]

OpenStudy (anonymous):

if it is not obvious, and my guess is it is not, use a calculator

OpenStudy (anonymous):

it is thanks so much :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

btw the answer is a whole number

OpenStudy (anonymous):

i know i got it right thanks

OpenStudy (anonymous):

yw